At the NASCAR race, the sound of the cars has a total power of 5,000 W. What decibel level would you hear 3 meters from the cars?


1
Expert's answer
2021-03-17T16:53:59-0400
"I=\frac{P}{4\pi r^2}=\frac{5000}{4\pi (3)^2}=44.2\frac{W}{m^2}"

"\beta=10\lg{\frac{44.2}{10^{-12}}}=136.45\ dB"